OpenStudy (anonymous):

Where does the graph of y = sin x from x = 0 to x = 2pi start?

jimthompson5910 (jim_thompson5910):

when x = 0, y is.. y = sin(x) y = sin(0) y = 0

jimthompson5910 (jim_thompson5910):

So the point (0,0) is on the graph of y = sin(x)

jimthompson5910 (jim_thompson5910):

When x = 2pi, y is.. y = sin(x) y = sin(2pi) y = 0 which means that the point (2pi, 0) is also a point on y = sin(x)